OpenStudy (anonymous):

What value for x could make the sentence true? x3 = 343

OpenStudy (anonymous):

1,029,49,21,7

OpenStudy (anonymous):

which option is correct

OpenStudy (anonymous):

is that 3 times x, or is that x^3? If it's x^3, then you need to cube root each side [\sqrt[3]{x ^{3}}=\sqrt[3]{343}\] to simplify/solve the equation. The answer would then be 7, as \[7\times7\times7=343\] which is to say \[7^{3}=343\]

OpenStudy (anonymous):

\[\sqrt[3]{x^{3}}=\sqrt[3]{343}\]